\begin{matrix}c \\ ◯ \end{matrix} R.J. van Glabbeek & C. Author This work is licensed under the Creative Commons Attribution License.

An Example of a Paper with a Rather Large Title-to-Content Ratio

Rob van Glabbeek

NICTA Sydney, Australia School of Computer Science and Engineering University of New South Wales^{∗} Sydney, Australia

                           rvg@cs.stanford.edu

Co Author

Independent Scholar

                                                                is@gmail.com

This is a sentence in the abstract. This is another sentence in the abstract. This is yet another sentence in the abstract. This is the final sentence in the abstract.

1 Introduction

The optional arguments of \documentclass{eptcs} are

   • at most one of draft, submission, preliminary or replacement,
   • at most one of publicdomain or copyright,
   • and optionally creativecommons,

– possibly augmented with

           ∗ noderivs
           ∗ or sharealike,

– and possibly augmented with noncommercial.

By means of the style-file option creativecommons authors equip their paper with a Creative Commons license that allows everyone to copy, distribute, display, and perform their copyrighted work and derivative works based upon it, but only if they give credit the way you request. By invoking the additional style-file option noderivs you let others copy, distribute, display, and perform only verbatim copies of your work, but not derivative works based upon it. Alternatively, the sharealike option allows others to distribute derivative works only under a license identical to the license that governs your work. Finally, you can invoke the option noncommercial that let others copy, distribute, display, and perform your work and derivative works based upon it but for noncommercial purposes only.

The correct values of \event, \volume, \anno and \firstpage will be communicated to you upon acceptance and should then be filled in in lines 2–5 of the the tex-file. Note that \event may contain an explicit newline command \\.

Authors’ (multiple) affiliations and emails use the commands \institute and \email. Both are optional. Authors should also supply \titlerunning and \authorrunning, using \def.

The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details [1].

EPTCS ??, 20??, pp. 1–3.

∗A fine university.

Exactly 46 lines fit on a page. 1 The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. 6 The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. 11 The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details.

Here starts a new paragraph. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. 16 The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. 21 The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. 26 The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. 31 The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. 36 The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. 41 The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. 46 The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details.

2 Bibliography

We request that you use \bibliographystyle{eptcs}. This style uses the bibtex fields ee or url, which are treated as synonyms, to make life links from your references to the response pages^{1} of the cited papers. We recommend only using this feature when you have archival-quality links: ones that promise to be valid for centuries. You can find archival-quality URL’s for most recently published papers in DBLP—they are in the bibtex-field ee. In fact, it is often useful to check your references against DBLP records anyway, or just find them there in the first place.

When using L^{A}TEX rather than pdflatex to typeset your paper, by default no linebreaking within long URLs is allowed. This leads often to very ugly output, that moreover is different from the output generated when using pdflatex. This problem is repaired when invoking \usepackage{breakurl}: it allows linebreaking within links and yield the same output as obtained by default with pdflatex. When invoking pdflatex, the package breakurl is ignored.

References

[1] R.J. van Glabbeek and C. Author (2008): An example of a paper with a rather large title-to-content ratio. Electronic Proceedings in Theoretical Computer Science 0, pp. 15–17. Available at http://style.eptcs.

   org/.

1Nowadays, papers that are published electronically tend to have a response page that lists the title, authors and abstract of the paper, and links to the actual manifestations of the paper (e.g. as dvi- or pdf-file). Sometimes publishers charge money to access the paper itself, but the response page is always freely available.